OpenStudy (anonymous):

The ratios in an equivalent ratio table are 3:12, 4:16, and 5:20. If the first number in the ratio is 10, what is the second number? Justify your reasoning.

OpenStudy (anonymous):

3:12 for example is 3*4=12

OpenStudy (anonymous):

So all of those the smallest number is being multiplied by 4

OpenStudy (anonymous):

So what do we do to 10?

OpenStudy (anonymous):

Multiply 10*4 and we will get 40!

OpenStudy (anonymous):

So the last ratio is 10:40 and all of them are proportional @maycie69

OpenStudy (anonymous):

Also to check 4*4=16 and 4*5=20

OpenStudy (anonymous):

thank you i have been tryen to solve this for 2 hours really thank u

OpenStudy (anonymous):

NO PROBLEM! :) fan and medal? @maycie69

OpenStudy (anonymous):

@maycie69 ??